# Market Making without Regret

### v1

@@ 3.1 $T^{\nicefrac{{2}}{{3}}}$ upper bound (Adversarial+Lip) or (IID+IV)
Then, for any $K\geq 2$ and any $K$-armed bandit algorithm $\mathcal{A}$, letting $R^{K}_{T}$ be the regret of $\mathcal{A}$ when the reward at any time $t\in[T]$ of any arm $k\in[K]$ is $(Z_{t}-q_{k})\mathbb{I}\{q_{k}\geq H_{t}\}$, the regret of Algorithm 3 run with parameters $K$ and $\mathcal{A}$ satisfies $R_{T}\leq R_{T}^{K}+\frac{\widetilde{L}+1}{2(K-1)}T\;,$ with $\widetilde{L}=L$ (resp., $\widetilde{L}=1$) if Item 1 (resp., Item 2) holds. In particular, if $T\geq 2$, by choosing $K\coloneqq\lceil T^{\nicefrac{{1}}{{3}}}\rceil+1$ and, as the underlying learning procedure $\mathcal{A}$, an adapted version of Poly INF [2], the regret of Algorithm 3 run with parameters $K$ and $\mathcal{A}$ satisfies $R_{T}\leq c\cdot T^{\nicefrac{{2}}{{3}}}\;,$ where $c\leq L+50$ (resp. $c\leq 51$) if Item 1 (resp., Item 2) holds.
A completely analogous theorem can be proved for Algorithm 4 for the problem of dynamic pricing with unknown costs (see Theorem B.1 in Appendix B).
input: Number of arms $K\in\{2,3,\dots\}$, Algorithm $\mathcal{A}$ for $K$-armed bandits
initialization: Let $q_{k}\coloneqq\frac{k-1}{K-1}$, for all $k\in[K]$
for time $t=1,2,\ldots$ do
      Let $I_{t}\in[K]$ be the output of $\mathcal{A}$ at time $t$;
      Post price $P_{t}\coloneqq q_{I_{t}}$ and observe feedback $C_{t}$ and $\mathbb{I}\{P_{t}<W_{t}\}$;
      Compute utility $(P_{t}-C_{t})\cdot\mathbb{I}\{P_{t}<W_{t}\}$ and feed it back to $\mathcal{A}$;
Algorithm 4 Discretized Bandits for Dynamic Pricing
Next, we use Theorem 3.2 and Theorem B.1 to show that the regret of M3 is bounded by $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$.
Theorem 3.3.
In the realistic-feedback online market-making problem, let $T\in\mathbb{N}$ be the time horizon and let $(M_{t},V_{t})_{t\in[T]}$ be the $[0,1]^{2}$-valued stochastic process representing the sequence of market prices and takers’ valuations. Assume that one of the two following conditions is satisfied:
@@
Then, let $K\coloneqq\lceil T^{\nicefrac{{1}}{{3}}}\rceil+1$, let $\mathcal{A}$ be the instance of Algorithm 3 that uses Poly INF as described in Theorem 3.2, and $\mathcal{A}^{\prime}$ be the instance of Algorithm 4 that uses Poly INF as described in Theorem B.1. Then the regret of M3 (Algorithm 2) run with parameters $\mathcal{A}$ and $\mathcal{A}^{\prime}$ satisfies $R_{T}\leq cT^{\nicefrac{{2}}{{3}}}$ with $c\leq 2L+100$ (resp., $c\leq 102$) if Item 1 (resp., Item 2) holds.
Proof.
@@
By Algorithm 2, the regret of M3 is upper bounded by the regret of a repeated first-price auctions problem with unknown valuations plus a dynamic pricing problem with unknown costs. Plugging in the bounds from Theorem 3.2 and Theorem B.1, we get the required result. ∎
3.2 $T^{\nicefrac{{2}}{{3}}}$ lower bound (IID+Lip+IV)
In this section we prove the main lower bound, showing that no algorithm can achieve a regret better than $\Omega(T^{\nicefrac{{2}}{{3}}})$ even when the sequence of market values and taker’s valuations is i.i.d. with a smooth distribution, and market values and taker’s valuations are independent of each other.
Theorem 3.4.
In the realistic-feedback online market-making problem, for each $L\geq 8$, each time horizon $T\geq 42$, and each (possibly randomized) algorithm for realistic-feedback online market making, there exists a $[0,1]^{2}$-valued i.i.d. sequence $(M_{t},V_{t})_{t\in[T]}$ of market values and taker’s valuations such that for each $t\in[T]$ the two random variables $M_{t}$ and $V_{t}$ are independent of each other, they admit an $L$-Lipschitz cumulative distribution function, and the regret of the algorithm over the sequence $(M_{t},V_{t})_{t\in[T]}$ is lower bounded by $R_{T}\geq cT^{\nicefrac{{2}}{{3}}}$ where $c$ is a constant and $c\geq 10^{-6}$.
Before presenting the full proof of this theorem, we give a high-level description of its key ideas. The first step is to build a base joint distribution over market values and takers’ valuations such that the expected utility function of the learner is maximized over an entire segment of pairs of prices $(b,1)$, for $b$ that belongs to some interval (which are instances where it is never optimal to sell). Then, we partition this set of maximizers into $K=\Theta(T^{\nicefrac{{1}}{{3}}})$ segments with the same size and build $K$ perturbations of the base distribution such that, in each one of these perturbations, the corresponding expected utilities have a small $\varepsilon=\Theta(1/K)$-spike inside one of these subsegments. To obtain this result, we draw market values uniformly on $\bigl{[}\frac{7}{8},1\bigr{]}$ and define a more involved distribution for the takers’ valuations (see Figure 5 for the plot of one of these perturbed distributions over takers’ valuations). From here, we start with a simple observation: If we content ourselves with proving a lower bound for algorithms that play exclusively bid/ask pairs of the form $(b,1)$, our problem reduces to repeated first-price auctions with unknown valuations. In this simplified problem, the learner has to essentially locate an $\varepsilon$-spike present in one of $\nicefrac{{1}}{{\varepsilon}}$ locations. Therefore, the learner can either refuse to locate the spike, consequently paying an overall $\Theta(\varepsilon T)$ regret in the worst case, or invest $\Omega(\nicefrac{{1}}{{\varepsilon^{2}}})$ rounds [^4] Given that the KL divergence between the base distribution of the feedback of points in a perturbed region and the perturbed distribution of the feedback of the same points is $O(\varepsilon^{2})$, a standard information-theoretic argument shows that $\Omega(\nicefrac{{1}}{{\varepsilon^{2}}})$ samples of points in a region are needed to determine if the spike is present in that region. in trying to locate each one of the $\nicefrac{{1}}{{\varepsilon}}$ potential spikes, paying $\Omega(\varepsilon)$ each time a point in the wrong region is selected, for an overall regret of $\Omega\bigl{(}\frac{1}{\varepsilon^{2}}\cdot\frac{1}{\varepsilon}\cdot\varepsilon\bigr{)}=\Omega(\nicefrac{{1}}{{\varepsilon^{2}}})$. The hardest instance is when $\varepsilon T=\nicefrac{{1}}{{\varepsilon^{2}}}$, i.e., when $\varepsilon=T^{\nicefrac{{1}}{{3}}}$. In this case, no matter what the learner decides, they will always pay at least $\Omega(\varepsilon T)=\Omega(\nicefrac{{1}}{{\varepsilon^{2}}})=\Omega(T^{\nicefrac{{2}}{{3}}})$ regret. The problem becomes substantially more delicate in our setting because, now, to explore a potential action $(b,1)$, the learner has access to uncountably more options $(b,a)$ (see Figure 3).
Figure 3: The heat maps of the amount of information (on the left) and the expected utilities (on the right) obtained by selecting pairs of prices in a hard instance we construct in our $\Omega(T^{\nicefrac{{2}}{{3}}})$ lower bound for i.i.d. smooth instances with market values independent of takers’ valuations (Theorem 3.4). To distinguish the optimal region with reward of (at least) $\nicefrac{{1}}{{8}}+c_{\mathrm{spike}}\varepsilon_{K}$ (the yellow region at the top of the right plot) from the suboptimal region with reward $\nicefrac{{1}}{{8}}$, one could play any point in the non-blue horizontal and vertical regions on the left plot. Note that the pairs that yield the highest amount of feedback (the two small yellow neighborhoods close to the diagonal on the left plot) give highly suboptimal rewards, which are at least $c_{\textrm{plat}}=\Omega(1)$ below $\nicefrac{{1}}{{8}}$.
This complicates things because, in usual online learning lower bounds, the worst-case regret of an algorithm is analyzed by counting how many times the algorithm plays exploiting and exploring arms, then quantifying how much information was gathered from the exploration and summing over all exploring arms. This strategy is hardly implementable in our setting because each exploiting pair $(b,1)$ can be explored with uncountably many other arms $(b^{\prime},s^{\prime})$, and each one trades off some reward to gather some amount of feedback (see Figure 4)—for example, one could post two prices $(b,b+\theta)$, for some small $\theta>0$ paying $\Omega(1)$ regret to obtain much higher quality feedback than if they played $(b,1)$ ($\varepsilon$ vs $\varepsilon^{2}$ KL-divergence information).
Figure 4: The amount of information (on the left) and the corresponding expected utility (on the right) as a function of pairs of prices $(b,a)$, with $(b,a)$ following the red dashed line in Figure 3 starting at the left boundary of the horizontal non-blue region at the point $(0,\widetilde{b})$, moving horizontally up to the diagonal, then vertically up to the optimal pairs $(b^{\star},1)$, parameterized by a real number $x$. The points $(\widetilde{b},\widetilde{a})$ and $(\widetilde{b},p_{\mathrm{exploit}})$ illustrate how regions of high reward and regions of high information do not coincide, in fact the cost of gathering more information (play $(\widetilde{b},\widetilde{a})$ instead of $(\widetilde{b},p_{\mathrm{exploit}})$) can be a constant $\Omega(1)$ in the reward. On the other hand, a high-reward play can cost $\Omega(\epsilon)$ in the amount of information (but still gathering a quantity $\Omega(\epsilon^{2})$).
We circumvent this problem by clustering the action set $\mathcal{U}$ into finitely many disjoint subsets (see Figure 6) and analyzing all points belonging to the same set in a similar way. This way, we are able to prove that no matter how an algorithm decides to play, there will always be instances where it has to pay a regret of order at least $\Omega(T^{\nicefrac{{2}}{{3}}})$.
Proof.
@@ 3.2 $T^{\nicefrac{{2}}{{3}}}$ lower bound (IID+Lip+IV)
We define a family of perturbations parameterized by the set $\Xi\coloneqq\{(r,\varepsilon)\in[\frac{3}{16},\frac{11}{16}]\times[0,1]\mid\frac{3}{16}\leq r-\frac{\varepsilon}{2}\leq r+\frac{\varepsilon}{2}\leq\frac{11}{16}\}$; for each $(r,\varepsilon)\in\Xi$, define $g_{r,\varepsilon}\coloneqq\frac{1}{9}\cdot\mathbb{I}_{[r-\frac{\varepsilon}{2},r]}-\frac{1}{9}\cdot\mathbb{I}_{(r,r+\frac{\varepsilon}{2}]}$ and $f_{r,\varepsilon}\coloneq f+g_{r,\varepsilon}$. Notice that for each $(r,\varepsilon)\in\Xi$ the function $f_{r,\varepsilon}$ is still a density function whose corresponding cumulative distribution function $F_{r,\varepsilon}$ satisfies, for each $x\in[0,1]$,
@@
$\left\lvert F_{r,\varepsilon}(a)-F_{r,\varepsilon}(b)\right\rvert=\int^{a}_{b}f_{r,\varepsilon}(x)\,\,\mathrm{d}x=\max_{c\in[b,a]}f_{r,\varepsilon}(c)(a-b)\leq\left(\max_{c^{\prime}\in[0,1]}f(c^{\prime})+\frac{1}{9}\right)(a-b)=4\,(a-b)\,,$

### v2

@@ 2 Setting
In the online market-making problem, for each $L\geq 8$, each time horizon $T\geq 42$, and each (possibly randomized) algorithm for online market making, there exists a $[0,1]^{2}$-valued i.i.d. sequence $(M_{t},V_{t})_{t\in[T]}$ of market prices and takers’ valuations such that for each $t\in[T]$ the two random variables $M_{t}$ and $V_{t}$ are independent of each other, they admit $L$-Lipschitz cumulative distribution functions, and the regret of the algorithm over the sequence $(M_{t},V_{t})_{t\in[T]}$ is lower bounded by $R_{T}\geq cT^{\nicefrac{{2}}{{3}}}$ where $c\geq 10^{-6}$ is a constant.
We present here a proof sketch and defer the full proof of this result to Appendix D.
Proof sketch
The first step is to build a base joint distribution over market prices and takers’ valuations such that the expected utility function of the learner is maximized over an entire segment of pairs of prices $(b,1)$, for $b$ that belongs to some interval—note that these are instances where it is never optimal to sell. Then, we partition this set of maximizers into $K$ segments with the same size and build $K$ perturbations of the base distribution such that, in each one of these perturbations, the corresponding expected utilities have a small $\varepsilon=\Theta(1/K)$-spike inside one of these subsegments.
We then make a simple observation: If we content ourselves with proving a lower bound for algorithms that play exclusively bid/ask pairs of the form $(b,1)$, our problem reduces to repeated first-price auctions with unknown valuations. In this simplified problem, the learner has to essentially locate an $\varepsilon$-spike among one of $\nicefrac{{1}}{{\varepsilon}}$ locations. Therefore, the learner can either refuse to locate the spike, and pay an overall $\Theta(\varepsilon T)$ regret, or invest $\Omega(\nicefrac{{1}}{{\varepsilon^{2}}})$ rounds [^6] Given that the KL divergence between the base distribution of the feedback of points in a perturbed region and the perturbed distribution of the feedback of the same points is $O(\varepsilon^{2})$, an information-theoretic argument shows that a sample of size $\Omega(\nicefrac{{1}}{{\varepsilon^{2}}})$ in any region is needed to determine whether the spike occurs in that region. in trying to locate each one of the $\nicefrac{{1}}{{\varepsilon}}$ potential spikes, paying $\Omega(\varepsilon)$ each time a point in the wrong region is sampled, for an overall regret of $\Omega\bigl{(}\frac{1}{\varepsilon^{2}}\cdot\frac{1}{\varepsilon}\cdot\varepsilon\bigr{)}=\Omega(\nicefrac{{1}}{{\varepsilon^{2}}})$. The hardest instance is when $\varepsilon T=\Theta(\nicefrac{{1}}{{\varepsilon^{2}}})$, i.e., when $\varepsilon=\Theta(T^{\nicefrac{{1}}{{3}}})$. In this case, no matter what the learner decides, they will always pay at least $\Omega(\varepsilon T)=\Omega(\nicefrac{{1}}{{\varepsilon^{2}}})=\Omega(T^{\nicefrac{{2}}{{3}}})$ regret.
However, the problem is substantially more delicate in the general case where, to explore a potentially optimal action $(b,1)$, the learner has access to uncountably many options $(b^{\prime},a^{\prime})$ (see Figure 2). This also leads to another complication. In standard online learning lower bounds, the worst-case regret of an algorithm is analyzed by counting how many times the algorithm plays exploring vs. exploiting arms, then quantifying how much information was gathered from the exploration, and finally summing over all exploring arms. This strategy is hardly implementable in our setting because each exploiting pair $(b,1)$ can be explored with uncountably many other arms $(b^{\prime},a^{\prime})$, where each arm trades off reward with feedback in a different way (see Figure 3)—for example, one could post two prices $(b,b+\theta)$, for some small $\theta>0$ paying $\Omega(1)$ regret to obtain much higher quality feedback than if they played $(b,1)$ ($\varepsilon$ vs. $\varepsilon^{2}$ bits of information).
Figure 3: The amount of information (on the left) and the corresponding expected utility (on the right) as a function of pairs of prices $(b,a)$, with $(b,a)$ following the red dotted line in Figure 2 starting at the left boundary of the horizontal non-blue region at the point $(0,\widetilde{b})$, moving horizontally up to the diagonal, then vertically up to the optimal pair $(\widetilde{b},1)$, parameterized by a real number $x$. The points $(\widetilde{b},\widetilde{a})$ and $(\widetilde{b},p_{\mathrm{exploit}})$ illustrate how regions of high reward and regions of high information do not coincide, in fact the cost of gathering more information (e.g., play $(\widetilde{b},\widetilde{a})$ instead of $(\widetilde{b},p_{\mathrm{exploit}})$) can be a constant $\Theta(1)$ in the reward. On the other hand, a high-reward play gathers only $\Theta(\varepsilon^{2})$ amount of information, hence costing $\Theta(\varepsilon)$ with respect to the maximum amount of information that can be gathered.
[…]
We conclude this section by remarking that, to the best of our knowledge, this is the first lower bound for a problem exhibiting such a continuum of trade-offs between reward and feedback. As our setting is arguably simple, we expect this phenomenon to occur in other applications of online learning to digital markets. Therefore, we believe the lower bound techniques developed in this work could be of independent interest.
5 Unlearnability without (lip) or (iv) Assumptions
In this section, we show that, even under the (iid) assumption, without smoothness (lip) or independence between market prices and takers’ valuations (iv), learning is impossible in general.
Theorem 5.1.
In the online market-making problem, for each time horizon $T$ and each $\delta\in(0,\nicefrac{{1}}{{2}})$ and each (possibly randomized) algorithm for online market making, there exists a $[0,1]^{2}$-valued i.i.d. sequence $(M_{t},V_{t})_{t\in[T]}$ of market prices and takers’ valuations such that the regret of the algorithm over the sequence $(M_{t},V_{t})_{t\in[T]}$ is lower bounded by $R_{T}\geq\left(\frac{1}{2}-\delta\right)\cdot T$.
We present here a proof sketch and defer the full proof of this result to Appendix E.
Proof sketch
[…]
6 Conclusions and Future Directions
We initiated an investigation of market making under an online learning framework and provided tight bounds on the regret under various natural assumptions. While the regret of various problems related to financial trading was previously investigated, most of these results are from the viewpoint of a liquidity seeker. Liquidity providers are crucial for the functioning of markets, giving practical impact to our theoretical findings.
Extending our analysis to full feedback, in which the learner observes the private valuations at the end of each round, would shed light on the cost of the more realistic feedback studied in this paper. In a different direction, we could consider market makers that could clear their position at a later time, which would make our setting more closely related to bandits-with-knapsacks (Badanidiyuru et al., 2018), which is known to be unsolvable in the adversarial setting (Immorlica et al., 2023).
[…]
NCB is partially supported by the EU Horizon CL4-2021-HUMAN-01 research and innovation action under grant agreement 101070617, project ELSA (European Lighthouse on Secure and Safe AI), by the MUR PRIN grant 2022EKNE5K (Learning in Markets and Society), and by the One Health Action Hub, University Task Force for the resilience of territorial ecosystems, funded by Università degli Studi di Milano (PSR 2021-GSA-Linea 6).
RC is partially supported by the FAIR (Future Artificial Intelligence Research) project, funded by the NextGenerationEU program within the PNRR-PE-AI scheme (M4C2, investment 1.3, line on Artificial Intelligence).
LF is partially supported by the EU Horizon CL4-2022-HUMAN-02 research and innovation action under grant agreement 101120237, project ELIAS (European Lighthouse of AI for Sustainability).
TC gratefully acknowledges the support of the University of Ottawa through grant GR002837 (Start-Up Funds) and that of the Natural Sciences and Engineering Research Council of Canada (NSERC) through grants RGPIN-2023-03688 (Discovery Grants Program) and DGECR2023-00208 (Discovery Grants Program, DGECR -- Discovery Launch Supplement).
Appendix A Notation
$T$ Time horizon
$K$ Step of the grid over $[0,1]$
$\mathcal{U}$ Upper triangle over $[0,1]^{2}$
Market making
$M_{t}$ Market price
$V_{t}$ Taker’s valuation
$B_{t}$ Bid presented by the learner
$A_{t}$ Ask presented by the learner
First-price auction with unknown valuations
$Z_{t}$ Unknown valuation of the auctioned item
$H_{t}$ Highest competing bid in the auction
$X_{t}$ Bid presented by the learner
Dynamic pricing with unknown costs
$C_{t}$ Unknown cost of the item for sale
$W_{t}$ Buyer’s valuation of the item
$P_{t}$ Price presented by the learner
Table 2: Notation
Appendix B Additional Related Works
There exists a vast literature that treats various trading-related tasks as specific stochastic control problems and solves them using techniques from stochastic control theory (Shreve, 2004, 2005; Cartea et al., 2015; Gatheral, 2011). However, most of that work assumes that the parameters of the underlying stochastic process are known or have been fitted to historical data in a previous “calibration” step. Fitting a distribution to data is challenging, especially in a market with hundreds of thousands of assets. Moreover, the distribution-fitting process is typically unaware of the downstream optimization problem the distribution is going to be fed into, and thus is unlikely to minimize generalization error on the downstream task. The field of online learning (Cesa-Bianchi and Lugosi, 2006; Shalev-Shwartz, 2012; Hazan, 2016; Orabona, 2019) provides adversarial or distribution-free approaches for solving exactly the kinds of sequential decision making problems that are common in financial trading, and even though some recent research does apply these approaches to trading, they are still far from being widely adopted. An early contribution to this area is Cover’s model of “universal portfolios” (Cover, 1991), where a problem of portfolio construction is solved in the case where asset returns are generated by an adversary. Cover showed that one can still achieve logarithmic regret with respect to the best “constantly rebalanced portfolio”, that ensures the fraction of wealth allocated to each asset remains constant through time. A long line of work has built up on his model to handle transaction costs (Blum and Kalai, 1997) and side information (Cover and Ordentlich, 1996), as well as optimizing the trade-off between regret and computation (Helmbold et al., 1998; Borodin et al., 2000, 2003; Zimmert et al., 2022). The problem of pricing derivatives such as options has also been tackled using online learning, where the prices are generated by an adversary as opposed to a geometric Brownian motion (DeMarzo et al., 2006; Abernethy et al., 2012, 2013). However, most of the problems considered in the online learning framework are from the point of view of a liquidity seeker as opposed to a liquidity provider, such as a market maker.
First-price auctions with unknown evaluations.
When restricted solely to the option of buying stock, our problem can be viewed as a series of repeated first-price auctions with unknown valuations. This scenario has received prior attention within the context of regret minimization (Cesa-Bianchi et al., 2024c; Achddou et al., 2021). The present work leverages these existing results but applies them within a more intricate setting. Cesa-Bianchi et al. (2024c) explored this problem with varying degrees of transparency, defined as the information revealed by the auctioneer, and provide a comprehensive characterization. In our framework, the level of transparency is higher than that encountered in the bandit case, orthogonal to the semi-transparent case, and less informative that the transparent case; this positioning offers a novel perspective on the problem. Achddou et al. (2021) investigated the repeated first-price auction problem within a fixed stochastic environment with independence assumptions and provide instance-dependent bounds, we tackle a similar problem in Lemma C.1 and present distribution-free guarantees. Other types of feedback have been considered, e.g., the case where the maximum bids $(V_{t})_{t\in[T]}$ form an i.i.d. process and are observed only when the auction is lost (Han et al., 2024), for the case where the private evaluations $(M_{t})_{t\in[T]}$ are i.i.d. it is possible to achieve regret $\widetilde{O}(\sqrt{T})$.
Dynamic pricing with unknown costs.
When considering solely the option of selling stock, our problem aligns with the well-established field of dynamic pricing with unknown costs. This area boasts a rich body of research, prior research analyzed the setting in which the learner has $I$ items to sell to $N$ independent buyers and, with regularity assumptions of the underlying distributions, can achieve a $O((I\log T)^{\nicefrac{{2}}{{3}}})$ regret bound against an offline benchmark with knowledge of the buyer’s distribution (Babaioff et al., 2015). It is known that in the stochastic setting and, under light assumptions on the reward function, achieve regret $O(\sqrt{T\log{T}})$ (Kleinberg and Leighton, 2003); this result has been expanded upon by considering discrete price distributions supported on a set of prices of unknown size $K$ and it has been shown to be possible to achieve regret of order $O(\sqrt{KT})$ (Cesa-Bianchi et al., 2019). Our work on dynamic pricing focuses on a scenario where the learner maintains an unlimited inventory and must incur an unknown cost per trade before realizing any profit.
Traditional finance.
Traditionally, the finance literature first fits the parameters of a stochastic process to the market and then optimizes trading based on those parameters. For example, the Nobel prize-winning Black-Scholes-Merton formula (Black and Scholes, 1973) assumes that the price of a stock follows a geometric Brownian motion with known volatility and prices an option on the stock by solving a Hamilton-Jacobi-Bellman equation to compute a dynamic trading strategy whose value at any time matches the value of the option. Many similar formulae have since been derived for more exotic derivatives and for more complex underlying stochastic processes—see (Hull, 2017) for an overview. A stochastic control approach has also been taken to solve the problem of “optimal trade execution,” which involves trading a large quantity of an asset over a specified period of time while minimizing market impact (Almgren and Chriss, 2001; Bertsimas and Lo, 1998), to compute a sequence of trades such that the total cost matches a pre-specified benchmark (Konishi, 2002), and also for the problem of market making (Cartea et al., 2015). Once again, the underlying stochastic process is assumed to be known. The two-step approach of first fitting a distribution and then optimizing a function over the fitted distribution is very popular even for simpler questions. For example, the problem of “portfolio optimization/construction” deals with computing the optimal allocation of your wealth into various assets in order to maximize some notion of future utility. The celebrated Kelly criterion (Shalunov et al., 2020; Thorp, 1975, 1969; Hakansson, 1975; MacLean et al., 2012) recommends one to use the allocation that maximizes the expected log return, and the mean-variance theory (Markowitz, 1952) (another Nobel prize-winning work) says one should maximize a sum of expected return and the variance of returns (scaled by some measure of your risk tolerance). However both assume that one knows the underlying distribution over returns.
Appendix C Missing details in the Proof of Theorem 3.1
To guide the reader through the proof of Theorem 3.1, we split it in several steps.
First, consider the problem of repeated first-price auctions with unknown valuations. Since the feedback received at the end of each round allows to compute the utility of the learner, a natural strategy is to simply discretize the interval $[0,1]$ and run a bandit algorithm on the discretization. The pseudo-code of this simple meta-algorithm can be found in Appendix C. The following result shows its guarantees.
{algorithm2e}
[…]
Lemma C.1.
@@ Appendix B Additional Related Works
Then, for any $K\geq 2$ and any $K$-armed bandit algorithm $\mathcal{A}$, letting $R^{K}_{T}$ be the regret of $\mathcal{A}$ when the reward at any time $t\in[T]$ of any arm $k\in[K]$ is $(Z_{t}-q_{k})\mathbb{I}\{q_{k}\geq H_{t}\}$, the regret of Appendix C run with parameters $K$ and $\mathcal{A}$ satisfies $R_{T}\leq R_{T}^{K}+\frac{\widetilde{L}+1}{2(K-1)}T\;,$ with $\widetilde{L}=L$ (resp., $\widetilde{L}=1$) if Item 1 (resp., Item 2) holds. In particular, if $T\geq 2$, by choosing $K\coloneqq\lceil T^{\nicefrac{{1}}{{3}}}\rceil+1$ and, as the underlying learning procedure $\mathcal{A}$, an adapted version of Poly INF (Audibert and Bubeck, 2010), the regret of Appendix C run with parameters $K$ and $\mathcal{A}$ satisfies $R_{T}\leq c\cdot T^{\nicefrac{{2}}{{3}}}\;,$ where $c\leq L+50$ (resp. $c\leq 51$) if Item 1 (resp., Item 2) holds.
Proof C.2.
The regret $R_{T}$ of Appendix C when the unknown valuations are $Z_{1},\dots,Z_{T}$ and the highest competing bids are $H_{1},\dots,H_{T}$ is
  1. 1.
    $\left\lvert\phi(q)-\phi(q^{\prime})\right\rvert=\Bigl{\lvert}\mathbb{E}\bigl{[}(Z_{t}-q)\mathbb{I}\{H_{t}\leq q\}-(Z_{t}-q^{\prime})\mathbb{I}\{H_{t}\leq q^{\prime}\}\bigr{]}\Bigr{\rvert}\\ \qquad\qquad=\Bigl{\lvert}\mathbb{E}\bigl{[}(Z_{t}-q)\mathbb{I}\{q^{\prime}<H_{t}\leq q\}+(q^{\prime}-q)\mathbb{I}\{H_{t}\leq q^{\prime}\}\bigr{]}\Bigr{\rvert}\\ \leq\bigl{(}\mathbb{P}[H_{t}\leq q]-\mathbb{P}[H_{t}\leq q^{\prime}]\bigr{)}+(q-q^{\prime})\leq(L+1)(q-q^{\prime}).$
  2. 2.
    $\delta_{K}\leq\frac{T}{K-1}\;.$
Define $\widetilde{L}\coloneqq L$ in the first case and $\widetilde{L}\coloneqq 1$ in the second case. Next, pick the Poly INF algorithm (Audibert and Bubeck, 2010) as the underlying learning procedure $\mathcal{A}$ and apply the appropriate rescaling of the utilities $x\mapsto\frac{x+1}{2}$, which is necessary because the utility yields values in $[-1,1]$, while Poly INF was designed for rewards in $[0,1]$, this costs a multiplicative factor of 2 in the regret guarantee. The regret $R_{T}$ of Appendix C can be upper bounded by
whenever $K=\lceil T^{\nicefrac{{1}}{{3}}}\rceil+1$ and the second inequality comes from the regret guarantees of the rescaled version of Poly INF (Audibert and Bubeck, 2010, Theorem 11).
[…]
Lemma C.3.
  1. 1.
  2. 2.
Then, for any $K\geq 2$ and any $K$-armed bandit algorithm $\mathcal{A}$, letting $R^{K}_{T}$ be the regret of $\mathcal{A}$ when the reward at any time $t\in[T]$ of any arm $k\in[K]$ is $(q_{k}-C_{t})\mathbb{I}\{q_{k}<W_{t}\}$, the regret of Appendix C run with parameters $K$ and $\mathcal{A}$ satisfies
$R_{T}\leq R_{T}^{K}+\frac{\widetilde{L}+1}{2(K-1)}T\;,$
with $\widetilde{L}=L$ (resp., $\widetilde{L}=1$) if Item 1 (resp., Item 2) holds. In particular, if $T\geq 2$, by choosing $K\coloneqq\lceil T^{\nicefrac{{1}}{{3}}}\rceil+1$ and, as the underlying learning procedure $\mathcal{A}$, an adapted version of Poly INF (Audibert and Bubeck, 2010), the regret of Appendix C run with parameters $K$ and $\mathcal{A}$ satisfies
$R_{T}\leq c\cdot T^{\nicefrac{{2}}{{3}}}\;,$
where $c\leq L+50$ (resp. $c\leq 51$) if Item 1 (resp., Item 2) holds.
Proof C.4.
As in Lemma C.1, we can bound $R_{T}$ for the two cases by controlling the discretization error $\delta_{K}$ on the grid via Lipschitzness (Item 1) or leveraging the structure of the expected utility around the maximum (Item 2). The bound on the regret $\mathcal{R}_{T}$ of $\mathcal{A}$ follows from the same guarantees on Poly INF (Audibert and Bubeck, 2010).
Next, we use Lemma C.1 and Lemma C.3 to show that the regret of an appropriately tuned M3 is bounded by $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$. This implies is the second part of Theorem 3.1, fully concluding its proof.
Lemma C.5.
In the online market-making problem, let $T\in\mathbb{N}$ be the time horizon and let $(M_{t},V_{t})_{t\in[T]}$ be the $[0,1]^{2}$-valued stochastic process representing the sequence of market prices and takers’ valuations. Assume that one of the two following conditions is satisfied:
@@
Then, let $K\coloneqq\lceil T^{\nicefrac{{1}}{{3}}}\rceil+1$, let $\mathcal{A}^{\mathrm{fpa}}$ be the instance of Appendix C that uses Poly INF as described in Lemma C.1, and $\mathcal{A}^{\mathrm{dp}}$ be the instance of Appendix C that uses Poly INF as described in Lemma C.3. Then the regret of M3 (Section 3) run with parameters $\mathcal{A}^{\mathrm{fpa}}$ and $\mathcal{A}^{\mathrm{dp}}$ satisfies $R_{T}\leq cT^{\nicefrac{{2}}{{3}}}$ with $c\leq 2L+100$ (resp., $c\leq 102$) if Item 1 (resp., Item 2) holds.
Proof C.6.
@@
By Section 3, the regret of M3 is upper bounded by the regret of a repeated first-price auctions problem with unknown valuations plus a dynamic pricing problem with unknown costs. Plugging in the bounds from Lemma C.1 and Lemma C.3, we get the required result.
Appendix D Proof of Theorem 4.1
The following is the proof of our $T^{2/3}$ lower bound (Theorem 4.1).
Proof D.1.
@@
We define a family of perturbations parameterized by the set $\Xi\coloneqq\{(r,\varepsilon)\in[\frac{3}{16},\frac{11}{16}]\times[0,1]\mid\frac{3}{16}\leq r-\frac{\varepsilon}{2}\leq r+\frac{\varepsilon}{2}\leq\frac{11}{16}\}$; for each $(r,\varepsilon)\in\Xi$, define $g_{r,\varepsilon}\coloneqq\frac{1}{9}\cdot\mathbb{I}_{[r-\frac{\varepsilon}{2},r]}-\frac{1}{9}\cdot\mathbb{I}_{(r,r+\frac{\varepsilon}{2}]}$ and $f_{r,\varepsilon}\coloneq f+g_{r,\varepsilon}$ (see Figure 4). Notice that for each $(r,\varepsilon)\in\Xi$ the function $f_{r,\varepsilon}$ is still a density function whose corresponding cumulative distribution function $F_{r,\varepsilon}$ satisfies, for each $x\in[0,1]$,
@@
$\left\lvert F_{r,\varepsilon}(a)-F_{r,\varepsilon}(b)\right\rvert=\int^{a}_{b}f_{r,\varepsilon}(x)\,\,\mathrm{d}x=\max_{c\in[b,a]}f_{r,\varepsilon}(c)(a-b)\\ \leq\left(\max_{c^{\prime}\in[0,1]}f(c^{\prime})+\frac{1}{9}\right)(a-b)=4(a-b)\,,$